OpenStudy (anonymous):

i do i put five square root of three X cubed minus two square root of twenty seven cubed into simplest radical form?

OpenStudy (mertsj):

\[5\sqrt{3x^3}-2\sqrt{27^3}\]

OpenStudy (mertsj):

\[5\sqrt{3x^2x}-2\sqrt{9^33^3}=5x \sqrt{3x}-2\sqrt{9^39^13^1}=5x \sqrt{3x}-2(81)\sqrt{3}\]

OpenStudy (mertsj):

You can finish it.